# Stimulus Modality Matters: Impact of Perceptual Evaluations from Different Modalities on Speech Emotion Recognition System Performance

### v1

@@ I Introduction
Prior studies trained Speech Emotion Recognition (SER) systems using labels elicited by the different types of stimuli. There are two main ways to obtain labels. One is giving raters audio-only emotional stimuli and letting them assign labels. For instance, the MSP-PODCAST emotion dataset [1] uses this scenario. The other is giving raters audio-visual emotional stimuli and letting them provide labels, and the IEMOCAP [2] emotion dataset is in the condition. While some papers have utilized labels derived from audio-visual stimuli [3] to train SER systems, others have focused on training SER models with labels elicited by audio-only stimuli [4]. This variation in methodology raises an intriguing open question: is the emotional labels elicited by multi-modal emotional stimuli different from training SER systems than those elicited by audio-only?
The emergence of speech self-supervised learning models (SSLMs) has significantly propelled advancements across a wide array of speech-related tasks [5], including SER. The current state-of-the-art frameworks in SER are primarily built upon these SSLMs [6]. To thoroughly investigate the research question concerning the advantage of utilizing labels elicited by multi-modal or single-modal emotional stimuli for training SER models, we conducted extensive experiments using SSLMs. Our approach involved training SER systems using labels elicited from various modalities, such as audio-only, facial-only, and audio-visual. This training was facilitated by leveraging the S3PRL toolkit [5], encompassing 14 SSLMs. We revealed notable differences in the performances of SER systems trained with the labels elicited by the mentioned modalities across three testing conditions defined by labels elicited from audio-only, facial-only, and audio-visual modalities.
@@ III Methodology
Most prior SER researches have mainly relied on annotations elicited by audio-visual stimulus as their learning objective [3], or the prior study did not specify labels elicited by which modalities they used [15]. However, the labels elicited by which modalities can lead to better performances have yet to be explored. To answer the question, we aim to compare the performances of SER systems trained with labels elicited by different modalities (e.g., face-only or audio-only) across various testing conditions according to modality stimulus.
Figure 1: The figure shows the multi-modal emotional stimulus in the CREMA-D emotion corpus.
@@ III-B Proposed All-inclusive Label Set and Rationales
We propose an innovative approach for training SER systems. This approach involves creating an all-inclusive label set (All) that integrates labels elicited from uni-modal and multi-modal emotional stimuli. By leveraging the comprehensive spectrum of emotional cues available from audio, visual, and audio-visual sources, we aim to harness the synergistic effect of multi-modal input to improve the performance of SER systems. This strategy is directly motivated by the natural human ability to perceive better and interpret emotions when more modal cues are available.
In the example of Table I, we summarize how to convert the raw annotators into the training/testing labels elicited by the different modalities. The all-inclusive label (All) considers all labels elicited by voice-only, facial-only, and audio-visual stimuli. It is important to note that the labels are distributional labels for the training stage, and they are converted into binary vectors when the values are higher than the defined threshold described in Section IV-C, which is 1/C, where C is the number of emotions. In this work, we have six emotions in total (threshold = 1/6), so the testing label of the audio-visual scenario (1,1,0,0,0,0) is different from others (1,1,0,0,0,1) after applying the threshold method introduced in [10]. We also follow [17] to allow the samples to have more than one emotion to reflect the nature of emotion perception that could involve mixed emotions from the psychology perspective [18].
@@ III-C SSLMs-based SER Framework
We adopt the SER models using 14 SSLMs as the backbone models by following the settings of the EMO-SUPERB [^2] https://github.com/EMOsuperb/EMO-SUPERB-submission [19] to train SER systems. We use SSLMs as they achieve SOTA results in SER. We leverage two mainstream categories of SSLMs, pre-trained using generative loss, DeCoAR 2 [20], Autoregressive Predictive Coding (APC) [21], VQ-APC [22], Non-autoregressive Predictive Coding (NPC) [23], TERA [24], and Mockingjay [25]), and discriminative loss (XLS-R-1B) [26], WavLM Large [27], Data2Vec-A [28], Hubert Large [29], wav2vec 2.0 Large (W2V2) [30], VQ wav2vec (VQ-W2V) [31], wav2vec (W2V) [32], wav2vec 2.0 Robustness (W2V2 R) [33] and Contrastive Predictive Coding (CPC) (M CPC)[34]). Additionally, we include the log mel filterbank (FBANK).
@@ IV-C Evaluation Metrics and Confidence Intervals
Specifically, a prediction for a particular class is deemed correct if its proportional representation among all predictions exceeds the threshold of ($1/C$), where $C$ is the total number of emotional courses under consideration. This strategic choice of threshold ensures that predictions are classified based on a fair representation criterion, aligning with methodologies previously described in the literature [10, 39]. Notice that we collect the predictions from each partition defined in the study [19] and then measure the performances in macro-F1 score with the average and the lower and upper bound of the confidence interval (CI) between 2.75% and 97.5% using the toolkit [40]. All results are single-run with the fixed random seed number.
@@ IV-D Models Training and Choice
We employ the AdamW optimizer [41] with a learning rate 0.0001. The batch size is set to 32, and the models are trained for 50 epochs. The best-performing models are selected based on the lowest loss value on the development set. All experiments use two Nvidia Tesla V100 GPUs with 32 GB of memory, requiring approximately 84 GPU hours. Our work is built upon the S3PRL [5] [^4] https://github.com/s3prl/s3prl, which is implemented using PyTorch [42] and HuggingFace library [43].
@@ V Results and Analysis
What is a correlation between labels elicited by various modalities? We employ the concordance correlation coefficient (CCC) [44] to assess the correlation between averaged labels under different conditions, as presented in Table II. Interestingly, the CCC between Voice and AV modalities is only 0.573, while the CCC between Face and AV is considerably higher at 0.805. Furthermore, the CCC between Voice and All (0.745) is lower than the CCC between Face and All (0.875), as well as the CCC between AV and All (0.913). The All modality exhibits overall higher correlations with other modalities, providing an additional justification for our proposal of the all-inclusive label set (All), which considers all labels from all modalities. This approach ensures a comprehensive consideration of the available information when analyzing emotion perception.
TABLE III: The table summarizes performances of the various SSLMS-based SER systems trained with the labels elicited by different modalities. The Face, Voice, and AV represent face-only, voice-only, and audio-visual, respectively. We use bold to represent the best performance according to each upstream model.
@@
TABLE IV: Overview of SER performances based on the WavLM Large in the cross-testing conditions. The AV represents “Audio-Visual”. For column, Overall, we also indicate the lower and upper bound of the confidence interval between 2.75% and 97.5% for each result (lower bound, upper bound) using [40] in macro-F1 scores. For other columns, we use F1-scores.
Train Set Test Set Overall (lower bound, upper bound) Angry Sad Disgust fear Neutral Happy
Voice Voice 0.7117 (0.7043,0.7185) 0.7738 0.6285 0.6301 0.6665 0.9206 0.6508
Face 0.6146 (0.6079,0.6215) 0.6979 0.5583 0.5804 0.5586 0.7488 0.5436
AV 0.6486 (0.6417,0.6548) 0.7394 0.6154 0.6115 0.6289 0.7173 0.5792
All 0.6873 (0.6803,0.6938) 0.7419 0.6428 0.6394 0.6242 0.8694 0.6059
Voice Face 0.5634 (0.5555,0.5703) 0.6139 0.5075 0.5233 0.4804 0.6713 0.5841
Face 0.6366 (0.6303,0.6433) 0.6597 0.5458 0.6061 0.5835 0.7414 0.6834
AV 0.6382 (0.6320,0.6445) 0.6559 0.5337 0.6052 0.5775 0.7415 0.7154
All 0.6406 (0.6340,0.6471) 0.6548 0.5572 0.6152 0.5662 0.7201 0.7301
Voice AV 0.6418 (0.6350,0.6483) 0.7164 0.6155 0.6100 0.5802 0.6949 0.6337
Face 0.6750 (0.6691,0.6812) 0.7484 0.5593 0.6507 0.6264 0.7628 0.7021
AV 0.7076 (0.7016,0.7138) 0.7650 0.6246 0.6705 0.6679 0.7742 0.7436
All 0.7085 (0.7025,0.7142) 0.7539 0.6398 0.6936 0.6442 0.7534 0.7662
What is the effect of stimulus modality on performances of SER systems? The results in Table III reveal intriguing differences in the SER model’s performance when trained on labels elicited from different modalities. For the evaluation annotations, Voice utilizes voice-only annotations, Face uses face-only annotations, and AV uses audio-visual annotations. The crucial insight in Table III is the performance variation when different modalities are elicited. The best overall macro-F1 score in 9 out of 15 experiments was achieved by models trained with the Voice, and 5 out of 15 experiments was achieved by models trained with the AV. These findings highlight the significant impact that the chosen annotation modality can have on the capability of SER systems to recognize emotions from speech accurately. Consequently, when developing SER models, it is crucial to carefully consider the modality used for annotating emotional labels, as this factor can substantially influence the model’s ability to capture the nuances of emotions in speech.
@@ VI Discussion and Limitations
While the study by Paulmann et al. [7] suggests that humans have enhanced emotion recognition with multimodal stimuli compared to single modalities, our research found that audio-only labels are the most effective for training SER systems when only speech input is available. Systems were evaluated using audio-only, visual-only, audio-visual, and all-inclusive labels, with the audio-only approach proving to be the most optimal. Multimodal emotion systems are not used since current SER systems predominantly rely exclusively on audio input. Moreover, to ensure that emotion predictions closely resemble human perceptions, speech-only emotion recognition evaluations must use labels derived from voice-only contexts. Additionally, the findings from the mentioned study might not apply to other emotion recognition systems, such as those based on facial expressions, text, or audio-visual inputs.

### v2

@@ I Introduction
Prior studies trained Speech Emotion Recognition (SER) systems using labels elicited by the different types of stimuli. There are two main ways to obtain labels. One is giving raters audio-only emotional stimuli and letting them assign labels. For instance, the MSP-PODCAST emotion dataset [1] uses this scenario. The other is giving raters audio-visual emotional stimuli and letting them provide labels, and the IEMOCAP [2] emotion dataset is in the condition. While some papers have utilized labels derived from audio-visual stimuli [3] to train SER systems, others have focused on training SER models with labels elicited by audio-only stimuli [4]. This variation in methodology raises an intriguing open question: are the emotional labels elicited by multi-modal emotional stimuli different from those used in training SER systems with audio-only inputs?
The emergence of speech self-supervised learning models (SSLMs) has significantly propelled advancements across a wide array of speech-related tasks [5], including SER. The current state-of-the-art frameworks in SER are primarily built upon these SSLMs [6]. To thoroughly investigate the research question concerning the advantage of utilizing labels elicited by multi-modal or single-modal emotional stimuli for training SER models, we conducted extensive experiments using SSLMs. In our approach, we trained SER systems using labels derived from various modalities, including audio-only, facial-only, and audio-visual inputs. This training utilized the S3PRL toolkit [5], which encompasses 14 self-supervised learning models (SSLMs). Our findings highlight significant differences in the performance of SER systems trained with labels elicited from these modalities when tested under three distinct conditions: audio-only, facial-only, and audio-visual label elicitation.
@@ III Methodology
Most prior SER researches have mainly relied on annotations elicited by audio-visual stimulus as their learning objective [3], or the prior study did not specify labels elicited by which modalities they used [15]. However, the relationship between the modalities from which labels are elicited and the resulting performance improvements has not yet been thoroughly investigated. To answer the question, we aim to compare the performances of SER systems trained with labels elicited by different modalities (e.g., face-only or audio-only) across various testing conditions according to modality stimulus.
Figure 1: The figure shows the multi-modal emotional stimulus in the CREMA-D emotion corpus and the proposed all-inclusive label.
@@ III-B Proposed All-inclusive Label Set and Rationales
We propose an innovative approach for training SER systems. This approach involves creating an all-inclusive label set (All) that integrates labels elicited from uni-modal and multi-modal emotional stimuli. By leveraging the comprehensive spectrum of emotional cues available from audio, visual, and audio-visual sources, we aim to harness the synergistic effect of multi-modal input to improve the performance of SER systems. This strategy is directly inspired by the inherent human capability to more accurately perceive and interpret emotions when multiple modal cues are available.
In the example of Table I, we summarize how to convert the raw annotations into the training/testing labels generated by the different modalities. The all-inclusive label (All) considers all labels elicited by voice-only, facial-only, and audio-visual stimuli. We consider a six-class emotion task, including anger (A), disgust (D), fear (F), happiness (H), sadness (S), and a neutral state (N). For voice-only elicitation, the annotations are {A*2, S*2, N*4}; for facial-only elicitation, the annotations are {A*4, S*2, N*2}; and for audio-visual elicitation, the annotations are {A*6, S*2}. The proposed all-inclusive label integrates all ratings, resulting in annotations of {A*12, S*6, N*6}. It is important to note that the labels are distributional labels for the training stage, and they are converted into binary vectors when the values are higher than the defined threshold described in Section IV-C, which is 1/C, where C is the number of emotions. In this work, we have six emotions in total (threshold = 1/6), so the testing label of the audio-visual scenario (1,1,0,0,0,0) is different from others (1,1,0,0,0,1) after applying the threshold method introduced in [10]. We follow [17] to allow the samples to have more than one emotion to reflect the nature of emotion perception that could involve mixed emotions from the psychology perspective [18].
@@ III-C SSLMs-based SER Framework
We adopt SER models using 14 SSLMs as the backbone models following the EMO-SUPERB settings [^2] https://github.com/EMOsuperb/EMO-SUPERB-submission [19] to train SER systems. We use SSLMs as they achieve SOTA results in SER. We leverage two mainstream categories of SSLMs, pre-trained using generative loss, DeCoAR 2 [20], Autoregressive Predictive Coding (APC) [21], VQ-APC [22], Non-autoregressive Predictive Coding (NPC) [23], TERA [24], and Mockingjay [25]), and discriminative loss (XLS-R-1B) [26], WavLM Large [27], Data2Vec-A [28], Hubert Large [29], wav2vec 2.0 Large (W2V2) [30], VQ wav2vec (VQ-W2V) [31], wav2vec (W2V) [32], wav2vec 2.0 Robustness (W2V2 R) [33] and Contrastive Predictive Coding (CPC) (M CPC)[34]). Additionally, we include the log mel filterbank (FBANK).
@@ IV-C Evaluation Metrics and Confidence Intervals
Specifically, a prediction for a particular class is deemed correct if its proportional representation among all predictions exceeds the threshold of ($1/C$), where $C$ is the total number of emotional courses under consideration. This strategic choice of threshold ensures that predictions are classified based on a fair representation criterion, aligning with methodologies previously described in the literature [10, 39]. Notice that we collect the predictions from each partition defined in the study [19] and then measure the performance in macro-F1 score with the average and lower and upper bound of the confidence interval (CI) between 2.75% and 97.5% using the toolkit [40]. All results are single-run with a fixed random seed number.
@@ IV-D Models Training and Choice
We employ the AdamW optimizer [41] with a learning rate of 0.0001. The batch size is set to 32, and the models are trained for 50 epochs. The best-performing models are selected based on the lowest loss value on the development set. All experiments use two Nvidia Tesla V100 GPUs with 32 GB of memory, requiring approximately 84 GPU hours. Our work is built upon the S3PRL [5] [^4] https://github.com/s3prl/s3prl, which is implemented using PyTorch [42] and HuggingFace library [43].
@@ V Results and Analysis
What is a correlation between labels elicited by various modalities? We employ the concordance correlation coefficient (CCC) [44] to assess the correlation between the averaged labels under different conditions, as presented in Table II. Interestingly, the CCC between Voice and AV modalities is only 0.573, while the CCC between Face and AV is considerably higher at 0.805. Furthermore, the CCC between Voice and All (0.745) is lower than the CCC between Face and All (0.875), as well as the CCC between AV and All (0.913). The All modality exhibits overall higher correlations with other modalities, providing an additional justification for our proposal of the all-inclusive label set (All), which considers all labels from all modalities. This approach ensures a comprehensive consideration of the information available when analyzing emotion perception.
TABLE III: The table summarizes performances of the various SSLMS-based SER systems trained with the labels elicited by different modalities. The Face, Voice, and AV represent face-only, voice-only, and audio-visual, respectively. We use bold to represent the best performance according to each upstream model. All values are in macro-F1 scores.
@@
TABLE IV: Overview of SER performances based on the WavLM Large in the cross-testing conditions. The AV represents “Audio-Visual”. For column, Overall, we also indicate the lower and upper bound of the confidence interval between 2.75% and 97.5% for each result (lower bound, upper bound) using [40] in macro-F1 scores. For other columns, we use sample-based
F1-scores. Train Set Test Set Overall (lower bound, upper bound) Angry Sad Disgust fear Neutral Happy Voice Voice 0.7117 (0.7043,0.7185) 0.7738 0.6285 0.6301 0.6665 0.9206 0.6508 Face 0.6146 (0.6079,0.6215) 0.6979 0.5583 0.5804 0.5586 0.7488 0.5436 AV 0.6486 (0.6417,0.6548) 0.7394 0.6154 0.6115 0.6289 0.7173 0.5792 All 0.6873 (0.6803,0.6938) 0.7419 0.6428 0.6394 0.6242 0.8694 0.6059 Voice Face 0.5634 (0.5555,0.5703) 0.6139 0.5075 0.5233 0.4804 0.6713 0.5841 Face 0.6366 (0.6303,0.6433) 0.6597 0.5458 0.6061 0.5835 0.7414 0.6834 AV 0.6382 (0.6320,0.6445) 0.6559 0.5337 0.6052 0.5775 0.7415 0.7154 All 0.6406 (0.6340,0.6471) 0.6548 0.5572 0.6152 0.5662 0.7201 0.7301 Voice AV 0.6418 (0.6350,0.6483) 0.7164 0.6155 0.6100 0.5802 0.6949 0.6337 Face 0.6750 (0.6691,0.6812) 0.7484 0.5593 0.6507 0.6264 0.7628 0.7021 AV 0.7076 (0.7016,0.7138) 0.7650 0.6246 0.6705 0.6679 0.7742 0.7436 All 0.7085 (0.7025,0.7142) 0.7539 0.6398 0.6936 0.6442 0.7534 0.7662
What is the effect of the stimulus modality on the performance of the SER systems? The results in Table III reveal intriguing differences in the performance of the SER model when trained on labels elicited from different modalities. For the evaluation annotations, Voice utilizes voice-only annotations, Face uses face-only annotations, and AV uses audio-visual annotations. The crucial insight in Table III is the performance variation when different modalities are elicited. The best overall macro-F1 score in 9 out of 15 experiments was achieved by models trained with the Voice, and 5 out of 15 experiments was achieved by models trained with the AV. These findings highlight the significant impact that the chosen annotation modality can have on the capability of SER systems to recognize emotions from speech accurately. Consequently, when developing SER models, it is crucial to carefully consider the modality used for annotating emotional labels, as this factor can substantially influence the model’s ability to capture the nuances of emotions in speech.
@@ VI Discussion and Limitations
While the study by Paulmann et al. [7] suggests that humans have enhanced emotion recognition with multimodal stimuli compared to single modalities, our research found that audio-only labels are the most effective for training SER systems when only speech input is available. Systems were evaluated using audio-only, visual-only, audio-visual, and all-inclusive labels, with the audio-only approach proving to be the most optimal. Multimodal emotion systems are not used since current SER systems predominantly rely exclusively on audio input. Moreover, to ensure that emotion predictions closely resemble human perceptions, speech-only emotion recognition evaluations must use labels derived from voice-only contexts. Additionally, the findings from the mentioned study might not apply to other emotion recognition systems, such as those based on facial expressions, text, or audio-visual inputs. Besides, our experiments have one limitation: they were conducted on a single emotion database, as no other publicly available databases provide emotional annotations across various stimuli.
@@ VII Conclusion and Future Work
Acknowledgment
This work was supported by the NSTC under Grant 113-2634-F-002-003. We thank Professor Hung-yi Lee for his valuable comments.